OpenStudy (jaylinnole19):

What volume of 0.716 M KBr solution is needed to provide 10.5 g of KBr

OpenStudy (master_333):

Use the formula for molarity to answer this question. Molarity = (moles of solute)/liters of solution First find the moles of the KBr n(# of moles) = mass/GMW n = 10.5g/ (39.1g + 79.9g) n = 0.088 moles Now place into the molarity equation Molarity = (moles of solute)/liters of solution 0.716 M = 0.088 moles/V V = 0.088/0.716 = 0.1229 L Therefore, the volume of 0.716 M KBr solution is needed to provide 10.5 g of KBr is 0.1229 L

OpenStudy (geobit):

\[n=\frac{ mass }{ molar mass }\] \[C=\frac{ no. moles }{ volume }\] hence finding vol. is \[V=\frac{ no. moles }{ conc.}\] so conc.=0.716M andmass=10.5g finding the moles you will get 0.088moles using the above mole formular. hence V=\[\frac{ 0.088 }{ 0.716 }=v\] V=0.1229L